### v1

@@ 5. Variance and Uncertainty Estimation:
[11] [6] formulas; Propagation to $\operatorname{Var}(S_{k})$ including OTU-level covariance; Alternatively, stratified bootstrap resampling.
@@ 6. Output:
As outlined in Figure 1, the workflow proceeds in six main steps: data preprocessing, log-contrast modeling, pairwise log-ratio screening, ranking and selection via the separability index $S_{k}$, uncertainty quantification, and final output generation. The following sections detail each of these steps.
@@ 4.1 Data description
We illustrate the proposed methodology using the colorectal cancer screening dataset originally reported by Baxter et al. [2] and later reanalyzed by Greenacre [10]. The data consist of operational taxonomic unit (OTU) abundances for 336 microbial features measured on stool samples from subjects classified as adenoma or control. In addition to the compositional microbiome data, the dataset includes several covariates, such as age, gender, and diabetes medication usage.
@@ 4.2 OTU relevance across models
The separation index estimates and their associated confidence intervals provide insight into the robustness and precision of OTU-based discrimination between adenoma and control groups. Model A, which does not adjust for any covariates, shows moderate separation performance, with a separation index around $0.58$. Adjustment for demographic covariates substantially improves discrimination: Model B, adjusting only for age, increases the separation index to approximately $0.67$, while Models C and D (adjusting for age and gender, and for age, gender, and diabetes medication, respectively) yiel
@@ 5 Discussion
Second, the influence of covariates on model performance and feature selection is evident. Accounting for variables such as age, gender, and diabetes medication improves model precision reflected in narrower confidence intervals for the separability index $S_{k}$ and reveals additional taxa that were not detected in unadjusted analyses. This result reinforces the importance of incorporating relevant covariates in microbiome studies to reduce confounding and to improve the interpretability of microbial associations with health outcomes.
@@ 6 Conclusions and future work
Looking ahead, several extensions are worth pursuing. One direction involves extending the framework to hierarchical taxonomic levels (e.g., genus or family), which could improve biological interpretation. Another promising line is the incorporation of Bayesian priors to stabilize feature selection in high-dimensional contexts. Scalability to other omics domains such as metagenomics, metabolomics, or transcriptomics will also broaden the applicability of the method. Finally, integrating cross-study validation procedures could help assess the generalizability and reproducibility of the identified microbial signatures.
@@
Declarations
  • Funding: R. Alberich, I. García Mosquera, A. Mir and F. Rosselló acknowledge support from the research project PID2021-126114NB-C44 funded by MCIN/AEI/ 10.13039/501100011033 and by “ERDF A way of making Europa”.

### v2

@@ 5. Variance and Uncertainty Estimation:
Hanley–McNeil ([8]) or DeLong ([5]) formulas; Propagation to $\operatorname{Var}(S_{k})$ including OTU-level covariance; Alternatively, stratified bootstrap resampling.
@@ 6. Output:
As outlined in Figure 1, the workflow proceeds in six main steps: data preprocessing, log-contrast modeling, pairwise log-ratio screening, ranking and selection via the separability index $S_{k}$, uncertainty quantification, and final output generation.
@@ 5.1 Data description
We illustrate the proposed methodology using the colorectal cancer screening dataset originally reported by [2] and later reanalyzed by [7]. The data consist of operational taxonomic unit (OTU) abundances for 336 microbial features measured on stool samples from subjects classified as adenoma or control. In addition to the compositional microbiome data, the dataset includes several covariates, such as age, gender, and diabetes medication usage.
@@ 5.2 OTU relevance across models
The separation index estimates and their associated confidence intervals provide insight into the robustness and precision of OTU-based discrimination between adenoma and control groups. Model A, which does not adjust for any covariates, shows moderate separation performance, with a separation index around $0.58$. Adjustment for demographic covariates substantially improves discrimination: Model B, adjusting only for age, increases the separation index to approximately $0.67$, while Models C and D (adjusting for age and gender, and for age, gender, and diabetes medication, respectively) yield separation indices close to 0.70, representing the strongest overall discrimination among the four models.
@@ 6 Discussion
Second, the influence of covariates on model performance and feature selection is evident. Accounting for variables such as age, gender, and diabetes medication improves model precision—reflected in narrower confidence intervals for the separability index $S_{k}$—and reveals additional taxa that were not detected in unadjusted analyses. This result reinforces the importance of incorporating relevant covariates in microbiome studies to reduce confounding and to improve the interpretability of microbial associations with health outcomes.
@@ 7 Conclusions and future work
Looking ahead, several extensions are worth pursuing. One direction involves extending the framework to hierarchical taxonomic levels (e.g., genus or family), which could improve biological interpretation. Another promising line is the incorporation of Bayesian priors to stabilize feature selection in ultra–high-dimensional contexts. Scalability to other omics domains—such as metagenomics, metabolomics, or transcriptomics—will also broaden the applicability of the method. Finally, integrating cross-study validation procedures could help assess the generalizability and reproducibility of the identified microbial signatures.